OpenStudy (anonymous):

CAN SOMEONE PPPLLLLZZZZ HELP!!!! what is the product of 2√9d*3d√5

OpenStudy (anonymous):

is it sqrt9d

OpenStudy (anonymous):

or d*sqrt9

OpenStudy (anonymous):

it is 2√9d*3d √5

OpenStudy (anonymous):

nvm i got the answer it was 18d√5d but can u help me with a diffrent one

OpenStudy (anonymous):

\[=2d \sqrt{3^{2}}3d \sqrt{5}\] =\[2(3)d*3d \sqrt{5}\] =\[18d ^{2}\sqrt{5}\]

OpenStudy (anonymous):

yes \

OpenStudy (anonymous):

what is the exact value of 14z^2/√z^9

OpenStudy (anonymous):

\[=14\sqrt{z ^{4}}/\sqrt{z ^{9}}\] =\[14/\sqrt{z ^{5}}\]

OpenStudy (anonymous):

thats not one of the options ill put the options up for u

OpenStudy (anonymous):

A. 14√z/z^7 B. 14√z/z^3 C.14√z/z^2 D.14/z^2

OpenStudy (anonymous):

ok splitting has to be changed \[14\sqrt{z}\sqrt{z ^{3}}/\sqrt{z ^{9}}\] =\[14\sqrt{z}/\sqrt{z ^{6}}\] =\[14\sqrt{z}/z ^{3}\]

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

ur welcome